OpenStudy (anonymous):

Tehere are 6 times as many union workers as there are non union workers at the acme plant.8 fewer than the number of union workers is equal to 4 times the number of non union workers.how many union workers are there at the plant?

OpenStudy (anonymous):

if union workers is x and nonunion is y try figuring the equations out

OpenStudy (anonymous):

can u explain how do i get the equation if u can please

OpenStudy (anonymous):

sure. 6union=nonunion so 6x=y 8-union=4*nonunion so, 8-x=4y solve the system

OpenStudy (anonymous):

okie

OpenStudy (anonymous):

try it out

OpenStudy (anonymous):

am trying

OpenStudy (anonymous):

ca not sorry

OpenStudy (anonymous):

yes cause i wrote it wrong sorry. the first is x=6y. try it out now. s

OpenStudy (anonymous):

still am still learning and i having trouble with this question

OpenStudy (asnaseer):

the equations have not been formed correctly. lets call the number of union workers "u" and the number of non-union workers "n" "There are 6 times as many union workers as there are non union workers" translates to: u = 6n "8 fewer than the number of union workers is equal to 4 times the number of non union workers" translates to: u - 8 = 4n so, the equations to solve are:\[u=6n\tag{1}\]\[u-8=4n\tag{2}\]substitute the expression for "u" from (1) into (2) and solve to find "n". then substitute this value of "n" into (1) to find "u".

OpenStudy (anonymous):

i feel puzzled now explion please i got 24 as an answer/

OpenStudy (asnaseer):

your answer is correct :)

OpenStudy (asnaseer):

which part is puzzling you?

OpenStudy (anonymous):

was

OpenStudy (anonymous):

ty

OpenStudy (asnaseer):

yw :)

OpenStudy (anonymous):

r u still around?